OpenStudy (anonymous):

How do you solve probabilities when it gives you two different spinners and ask you to find P(pink and T) or P (Blue an not X)?

OpenStudy (anonymous):

probability is simply amt of what you are looking for divided by total. so if one spinner has 8 equal sections and one is pink, the probability for that spinner to hit pink is 1/8. without seeing pictures for this question in particular, I can not give you exact details.

OpenStudy (anonymous):

One spinner has 4 equal sections: Red Blue Red Pink The second spinner has 8 equal parts T T H H X T T X . I understand the probability of pink is 1/4 and probabilties of T is 4/8. Am I suppose to combine them both since the question is P(pink and T)

OpenStudy (anonymous):

when it has to be both, multiply them together.

OpenStudy (anonymous):

All together like the 1/4*4/8=4/32?? Does that sound right?

OpenStudy (anonymous):

yup then simplify it. 4/32=1/8

OpenStudy (anonymous):

so blue and not x. 1/4 and 6/8=2/3 1/6

OpenStudy (anonymous):

Yes! Thanks I am slowing getting this.

OpenStudy (anonymous):

yay thats what this place is about. good luck.

OpenStudy (anonymous):

Do you know if a probability has or like...P (A or B) in it what you are suppose to do.